OpenStudy (anonymous):

Find the product. -2x(x 2 - 3) -2x3 - 6x -2x3 + 6x -2x2 + 6x 6x3

OpenStudy (anonymous):

\(\large (-2x)\times (x^2-3)=(-2x)\times x^2+(-2x)\times(-3)=-2x^3+6x\) The answer is C